OpenStudy (anonymous):

The circumference of a circle is 50.24 centimeters. Find the radius of the circle. (Use 3.14 for π.) 8 cm 16 cm 4 cm 2 cm

OpenStudy (anonymous):

\[C = 2 \pi r\]\[50.24 \approx 2 \times 3.14 \times r\] Solve for r.

OpenStudy (anonymous):

2X3.14=6.28

OpenStudy (anonymous):

\[r = \frac{50.24}{2 \times 3.14}\]

OpenStudy (anonymous):

do u divide

OpenStudy (anonymous):

@ghuczek

OpenStudy (anonymous):

its 8